# Simultaneous Reward Distillation and Preference Learning: Get You a Language Model Who Can Do Both

### v1

@@ 4 Direct Reward Distillation and policy-Optimization (DRDO)
The first step in DRDO involves training an Oracle model $\mathcal{O}$ to act as a robust proxy of true human preference inferred from the preference data. Second, we use $\mathcal{O}$ as a teacher to align the policy model (student) with a knowledge-distillation-based multi-task loss (Hinton, 2015; Gou et al., 2021) that regresses the student’s rewards onto those assigned by the Oracle. Within this multi-task setting, the student model simultaneously draws additional supervision from binary preference labels to make the most efficient use of finite preference data.
@@
Crucially, since the student model’s alignment depends on the quality of the Oracle and its generalizability, we use Yang et al. (2024)’s strategy to optimize $\mathcal{O}$ while retaining its language generation abilities. The core idea here is that regularizing the shared hidden states with a language generation loss in addition to the traditional RLHF-based reward modeling (Eq. 2) improves generalization ability for out-of-distribution preferences. This also helps minimize reward hacking (Kumar et al., 2022; Eisenstein et al., 2024), especially in offline settings. As such, our Oracle is optimized to minimize the following objective:
@@
$\mathcal{L}_{\mathcal{O}}(r_{\phi},\mathcal{D}_{\text{pref}})=-\mathbb{E}_{(x,y_{w},y_{l})\sim\mathcal{D}_{\text{pref}}}\left[(1-\alpha)(\log\sigma(r_{\phi}(x,y_{w})-r_{\phi}(x,y_{l})))+\alpha\log(y_{w})\right]$ (5)
@@
where $\alpha$ is the strength of language-generation regularization and $\phi$ represents the parameters of $\mathcal{O}$ being estimated. For more details on hyperparameter $\alpha$, see Section 5.
@@
A converged Oracle $\mathcal{O}$ can plausibly estimate true pointwise reward differences $r^{*}(x,y_{1})-r^{*}(x,y_{2})$ for any unlabeled sample $(x,y_{1},y_{2})$, without needing explicit access to preference labels. To fully leverage the diversity of preference signals, including non-deterministic preferences, we formulate the DRDO loss to utilize these signals directly. Specifically, the student model $\pi_{\theta}$ is optimized to match the Oracle’s reward differences using its own predictions, aligning closely with the Oracle’s behavior, while $\mathcal{O}$ itself does not get updated. This is achieved by minimizing a loss that combines both a supervised $\ell^{2}$-norm term and a novel focal-softened (Lin et al., 2018; Welleck et al., 2020) log odds-unlikelihood component:
@@
where $p_{w}=\sigma(z_{w}-z_{l})=\frac{1}{1+e^{-(z_{w}-z_{l})}}$ and quantifies the student policy’s confidence in correctly assigning the preference from $z_{w}=\log\pi_{\theta}(y_{w}\mid x)$ and $z_{l}=\log\pi_{\theta}(y_{l}\mid x)$, or the log-probabilities of the winning and losing responses, respectively. $\gamma$ and $\alpha$ are hyperparameters that regulate the strength of the modulating term and weighting factor [^4] Note that we do not use $\alpha$ as it is traditionally used in focal loss for weighting class imbalances (Mukhoti et al., 2020). Instead, since the true preference distribution is unknown, we tune the empirical optimal value based on validation data and keep it fixed during training. respectively.
@@
Intuitively, DRDO’s first component in the loss effectively regresses over the Oracle’s reward differences across all response pairs $(x,y_{1},y_{2})$ in $\mathcal{D}_{\text{pref}}$, using $\pi_{\theta}$’s own reward estimates, $\hat{r}_{1}$ and $\hat{r}_{2}$. Assuming the Oracle’s reward estimates correctly reflect the true preference strength, this simple distillation component is precisely minimized when the student $\pi_{\theta}$ perfectly emulates the Oracle’s reward difference assignment, without requiring it to explicitly match the pointwise rewards. The square over this term ensures that training is stabilized especially since we allow this component to fully regularize the loss without applying any weighting terms.
@@
For the preference component in DRDO loss (contrastive log-unlikelihood in Eq. 4), one can immediately draw a comparison with DPO which uses a fixed $\beta$ parameter (Eq. 3). DRDO uses a modulating focal-softened term, $(1-p_{w})^{\gamma}$ where $\pi_{\theta}$ learns from both deterministic and non-deterministic preferences, effectively blending reward alignment with preference signals to guide optimization. Intuitively, unlike DPO’s fixed $\beta$ that is applied across the whole training dataset, this modulating term amplifies gradient updates when preference signals are weak ($p_{w}\approx 0.5$) and tempering updates when they are strong ($p_{w}\approx 1$), thus ensuring robust learning across varying preference scenarios. When $\pi_{\theta}$ assigns high confidence to the winning response ($p_{w}\approx 1$, see Eq. 4), the focal loss contribution diminishes, reflecting minimal penalty due to strong deterministic preference signals. However, for harder cases with non-deterministic true preference ($p^{*}(y\succ y^{\prime})=(p^{*}(y^{\prime}\succ y)$), the focal term $\alpha(1-p_{w})^{\gamma}$ keeps DRDO gradient updates active and promotes learning even when preferences are ambiguous (Figure 2). This adaptive behavior ensures that DRDO maintains effective preference learning across varying preference strengths, especially when conventional methods like DPO struggle. See Lemma 1 and Lemma 2 for in-depth analysis. As shown in the gradient analysis in Appendix A with empirical proof in Fig. 4 (bottom-right), this modulating term acts like DPO’s gradient scaling term, in that it scales the DRDO gradient when the model incorrectly assigns preferences to easier samples.
@@ 5 Experimental Setup
Our experimental setup addresses two questions: How robust is DRDO alignment to nuanced or diverse preferences, in OOD settings?, and How well does DRDO achieve reward distillation with respect to model size? We empirically investigate these questions on two tasks: summarization and single-turn instruction following. Our experiments, including choice of datasets and models for each task are designed to to validate our approach as robustly as possible, subject to research budget constraints (see Section B.1 for more). We compare our approach with competitive baselines such as DPO (Rafailov et al., 2024) and e-DPO (Fisch et al., 2024), including the supervised finetuned (baseline) versions depending on the experiment. Some minor notes on the experimental setup described below can be found in Appendix B.
@@
We evaluate this on the Reddit TL;DR summarization dataset (Völske et al., 2017; Stiennon et al., 2020). We split the training data—Reddit forum posts—based on human labeler confidence in their preference annotation (high-confidence vs. low-confidence) and token edit distance between the preferred and the dispreferred responses (high-edit distance vs. low-edit distance). This results in two splits: $\mathcal{D}_{hc,he}$ and $\mathcal{D}_{\ell c,\ell e}$. Each contains $\sim$10k training samples, where the former comprises samples from the upper 50th percentile of the confidence and edit distance scores, mutatis mutandis. Intuitively, $\mathcal{D}_{hc,he}$ and $\mathcal{D}_{\ell c,\ell e}$ represent “easy” (deterministic) and “hard” (non-deterministic) preference samples where combined labeler confidence and string-dissimilarity act as proxy for the extreme ends of preference strengths/signals. See Appendix B.2 for more details. We denote the original training data $\mathcal{D}_{all}$. For each split, we compare our approach to DPO and e-DPO, where all baselines are initialized with Phi-3-Mini-4K-Instruct weights (Abdin et al., 2024) with supervised fine-tuning (SFT) on Reddit TL;DR human-written summaries. The test set is composed of news articles from CNN Daily, and therefore an out-of-distribution (OOD) setting.
@@
We evaluate all baselines on the cleaned version of the Ultrafeedback dataset (Cui et al., 2024) This experiment is conducted on the OPT suite of models (Zhang et al., 2022), at 125M, 350M, 1.3B, and 2.7B parameter sizes. The student policy is trained with SFT on the chosen responses of the dataset, following Rafailov et al. (2024). We exclude larger OPT models to focus on testing our distillation strategy with full-scale training, rather than parameter-efficient methods (PEFT) (Houlsby et al., 2019; Hu et al., 2021) to allow a full-fledged comparison considering all trainable parameters of the base model. For completeness and comparison across model families, we also include Phi-3-Mini-4K-Instruct following the same initialization.
DRDO and e-DPO Specifics
@@
TL;DR provides human-written reference summaries, so we use a high-capacity Judge to compute win-rates against baselines on 1000 randomly-chosen samples from the TL;DR test set. Following Rafailov et al. (2024), we use GPT-4o to compare the conciseness and the quality of the DRDO summaries and baseline summaries, while grounding its ratings to the human written summary. See Appendix E for our prompt format. For instruction following on Ultrafeedback, we sample generations from DRDO and all baselines at various diversity-sampling temperatures and report win-rates on the Ultrafeedback test set against $\mathcal{O}$. Following Lambert et al. (2024), we consider a win to be when, for two generations $y_{1}$ and $y_{2}$, we get $r(x,y_{1})>r(x,y_{2})$, where $r(x,y_{1})$ and $r(x,y_{2})$ are the expected rewards (logits) from the policies being compared. We also evaluate DRDO (trained on Ultrafeedback) against baselines (all trained on Ultrafeedback) on AlpacaEval, another instruction following benchmark for which we evaluating using GPT-4 Turbo as suggested by the creators (Li et al., 2023). Hyperparameters and model configurations are given in Appendix D.
@@ 6 Results
Table 1 shows the win rates computed with GPT-4o as judge for 1,000 randomly selected sample-prompts $x$ from the CNN Daily TL;DR evaluation set (Nallapati et al., 2016) under OOD settings. We follow similar settings as Rafailov et al. (2023) but further ground the prompt using human-written summaries as reference for GPT to conduct its evaluation (see Appendix E). For a fairer evaluation (Wang et al., 2024b; Goyal et al., 2023; Rafailov et al., 2023), we swap positions of $\pi_{\theta}$-generated summaries $y$ in the prompt to eliminate any positional bias and evaluate the generated summaries on criteria like coherence, preciseness and conciseness, with the human written summaries explicitly in the prompt to guide evaluation. Note that the win rate does not necessarily equal $N/1000$ because for a small number of samples ($<$ 10) GPT-4o did not return a judgment. These samples were discarded from the denominator.
Table 1: Win rates computed with GPT-4o as judge for 1,000 randomly selected samples from the evaluation set of the TL;DR CNN Daily dataset under OOD settings. This comparison evaluates the model against a baseline, presenting win counts alongside win rates. $\mathcal{D}_{all}$, $\mathcal{D}_{hc,he}$, and $\mathcal{D}_{\ell c,\ell e}$ represent TL;DR training splits. “Gold” samples refer to human-written reference summaries used in prompts to ground the win rate computations.
Comparison # Wins Win Rate
@@
For all policies $\pi_{\theta}$ trained on all three splits of the training data: $\mathcal{D}_{all}$, $\mathcal{D}_{hc,he}$, and $\mathcal{D}_{\ell c,\ell e}$, we compute the win-rates of DRDO vs. e-DPO and DPO to evaluate how each method performs at various levels of preference types. Across all settings, DRDO policies significantly outperform the two baselines. For instance, DRDO’s average win rates are almost 79.4% and 79.9% against e-DPO and DPO, respectively. As we had hypothesized, DRDO-aligned $\pi_{\theta}$ is able to learn preferences more effectively, especially in $\mathcal{D}_{\ell c,\ell e}$ the subset containing more non-deterministic preference samples. This suggests that DRDO is more robust to OOD-settings at various levels of difficulty in learning human preferences.
Reward distillation
Table 2 shows results from our evaluation of DRDO’s reward model distillation framework compared to DPO and e-DPO as well as baseline SFT-trained policies on the Ultrafeedback evaluation data, when compared across various model parameter sizes and at varying levels of temperature sampling. We sample $\pi_{\theta}$-generated responses to instruction-prompts in the test set using top-p (nucleus) sampling (Holtzman et al., 2019) of 0.8 at various temperatures $\in\{0.2,0.5,0.7,0.9\}$.
Table 2: Average win-rates computed with DRDO’s Oracle reward model against all baselines—SFT, DPO and e-DPO—at various diversity sampling temperatures ($T$). Bolded numbers show the highest win-rates against each baseline at each temperature.
Baselines
$T=0.2$ $T=0.5$ $T=0.7$ $T=0.9$
Policy (Student) SFT DPO e-DPO SFT DPO e-DPO SFT DPO e-DPO SFT DPO e-DPO
OPT-125M 0.47 0.61 0.51 0.52 0.63 0.50 0.57 0.58 0.54 0.57 0.64 0.61
OPT-350M 0.58 0.63 0.60 0.61 0.70 0.65 0.60 0.70 0.69 0.66 0.70 0.68
OPT-1.3B 0.83 0.70 0.83 0.82 0.65 0.81 0.83 0.63 0.75 0.81 0.59 0.78
OPT-2.7B 0.81 0.68 0.83 0.80 0.64 0.80 0.79 0.62 0.78 0.79 0.60 0.75
Phi-3 0.66 0.80 0.57 0.71 0.83 0.58 0.75 0.86 0.60 0.76 0.88 0.66
DRDO significantly outperforms competing baselines, especially for larger models in the OPT family. DRDO-trained OPT-1.3B, OPT-2.7B, and Phi-3-Mini-4K-Instruct achieve average win rates of 76%, 74%, and 72%, respectively, across all baselines. This is notable as responses are sampled on unseen prompts, and DRDO’s policy alignment is reference-free. DRDO’s robustness to diversity sampling further boosts performance, up to an 88% win rate against DPO with the Phi-3-Mini-4K-Instruct model. At lower temperatures, DRDO’s posted gains are more modest. Our results also indicate that performance is correlated with model size, as DRDO policies of the same size as the Oracle (1.3B) show the strongest gains. In smaller models, results are more mixed. [^6] This may be due to spurious length-correlation (Singhal et al., 2024; Meng et al., 2024), as smaller models generated avg. $\sim$211.9 tokens compared to their policy counterparts’ $\sim$190.8 tokens, which is closer to the original preferred responses in Ultrafeedback ($\sim$168.8). See Appendix D for more details.. DRDO shows moderate improvement and posts smaller gains against SFT models.
Table 3: Win rates on the AlpacaEval benchmark with GPT-4 Turbo as judge ($N=805$).
Comparison # Wins Win Rate
@@
Next, to precisely pinpoint DRDO’s instruction following ability in OOD settings, we further evaluate on the popular automatic evaluation benchmark AlpacaEval (Li et al., 2023), where LLM responses are specifically evaluated on 805 instructions chosen to be representative of how human-preferred policies would respond. Table 3 shows that, within controlled evaluation settings, GPT 4-Turbo consistently prefers DRDO responses 62.03% and 61.61% times compared to e-DPO and DPO respectively.
7 Analysis
Table 4 shows example generations from DRDO and a competitor where the DRDO example was preferred by the automatic judge. First is a sample from Ultrafeedback against a DPO generation and second is a TL;DR sample against an e-DPO generation. In both cases, we can see that the DRDO response is more concise and on-topic and the competitor output was dispreferred for reasons like condescending to the user, or including extraneous text about fulfilling the request.
@@ 7 Analysis
Using GPT-4o as a judge to approximate true human preferences may be prone to bias and there could be a gap, so we further validate our approach by investigating Oracle reward advantage over the above mentioned human written summaries as well as on-policy generations from an SFT-trained model. Figure 3 shows the computed expected reward advantages on the CNN Daily TL;DR evaluation set, sampled according to the method outlined in Section 6. Rewards were then computed using our Oracle (trained with Phi-3) on these sampled generations and normalized. To compute the advantage, we used SFT-trained model summaries (Figure 3[L]) and human written summaries (Figure 3[R]). While the expected reward advantage over human-written summaries is less than those with the baseline SFT policy, DRDO results in performance improvement across various temperature samplings. DRDO alignment appears in-tune with improving over a baseline SFT policy as our results on instruction tuning also suggest, brings in a considerable performance gain over competitive baselines like DPO and ensemble-based e-DPO while also being robust to OOD settings.
Figure 3: Oracle expected reward advantage on CNN Daily articles.
Our evaluations vs. “gold” summaries in Table 1 also demonstrate where bias may arise in the GPT-4o evaluation. The last three rows compare the model’s performance directly against these gold summaries (win counts and win rates are shown for gold samples). GPT-4o narrowly prefers generated summaries to human-written ones. One possible reason could be that Reddit TL;DR data is massive and crowd-sourced which naturally results in noisy labels. However, under this experiment too, DRDO-trained policies are better-performing than e-DPO and DPO, by about 1.5–6%. We should note that human summaries may contain more implicit diversity than generated summaries, and this may demonstrate the “regression to the mean” effect in LLM generation (Wu et al., 2024).

### v2

@@ 4 Direct Reward Distillation and policy-Optimization (DRDO)
DRDO alignment involves two steps. First, we train an Oracle model $\mathcal{O}$ to act as a robust proxy of true human preference inferred from the preference data. Second, we use $\mathcal{O}$ as a teacher to align the policy model (student) with a knowledge-distillation-based multi-task loss (Hinton, 2015; Gou et al., 2021) that regresses the student’s rewards onto those assigned by the Oracle. The student model simultaneously draws additional supervision from binary preference labels to make the most efficient use of finite data.
@@
Crucially, since the student model’s alignment depends on the quality of the Oracle and its generalizability, we use Yang et al. (2024)’s strategy to optimize $\mathcal{O}$ while retaining its language generation abilities. The core idea here is that regularizing the shared hidden states with a language generation loss in addition to the traditional RLHF-based reward modeling (Eq. 2) improves generalization to out-of-distribution preferences. This is achieved by initializing a separate linear reward head (parametrized by $\phi^{\prime}$) [^2] For simplicity of notation, on the LHS of Eq. 4 we subsume parameters $\phi^{\prime}$ into $\phi$. on top of the base LM (parametrized by $\phi$), which adds a negligible 0.003% additional parameters compared to the LM’s language modeling head. This also helps minimize reward hacking (Kumar et al., 2022; Eisenstein et al., 2024), especially in offline settings. As such, $\mathcal{O}$ is optimized to minimize the following objective:
@@
$\begin{split}\mathcal{L}_{\mathcal{O}}(r_{\phi},\mathcal{D}_{\text{pref}})&=-\mathbb{E}_{(x,y_{w},y_{l})\sim\mathcal{D}_{\text{pref}}}\\ &\left[(1-\alpha)(\log\sigma(r_{\phi^{\prime}}(x,y_{w})-r_{\phi^{\prime}}(x,y_{l})))+\alpha\log(r_{\phi}(y_{w}))\right]\end{split}$
@@
where $\alpha$ is the strength of language-generation regularization on the winning response log-likelihoods assigned by $\mathcal{O}$ (denoted $\log(r_{\phi}(y_{w})$) and $\phi$ represents the parameters of $\mathcal{O}$ being estimated.
Student Policy Alignment to Rewards and Preferences
@@
A converged Oracle $\mathcal{O}$ can plausibly estimate true pointwise reward differences $r^{*}(x,y_{1})-r^{*}(x,y_{2})$ for any unlabeled sample $(x,y_{1},y_{2})$, without needing explicit access to preference labels. We formulate the DRDO loss to directly leverage the diversity of preference signals, including non-deterministic preferences. Specifically, the student model $\pi_{\theta}$ is optimized to match its own reward predictions $\hat{r_{1}}$ and $\hat{r_{2}}$ to $\mathcal{O}$’s reward differences, aligning closely with $\mathcal{O}$’s behavior, while $\mathcal{O}$ itself does not get updated. The student model’s reward estimates are computed with a linear reward head initialized on top of the base LM, similar to Oracle training. This optimization is achieved by a knowledge-distillation loss ($\mathcal{L}_{\mathrm{kd}}$) that combines both a supervised $\ell^{2}$-norm term and a novel focal-softened (Lin et al., 2018; Welleck et al., 2020) log odds-unlikelihood component:
$\mathcal{L}_{\mathrm{kd}}(r^{*},\pi_{\theta})=\mathbb{E}_{(x,y_{1},y_{2})\sim\mathcal{D}_{\text{pref}}}\Bigg{[}\\ \underbrace{\left(r^{*}(x,y_{1})-r^{*}(x,y_{2})-(\hat{r}_{1}-\hat{r}_{2})\right)^{2}}_{\text{Reward Difference}}\quad-\\ \underbrace{\alpha(1-p_{w})^{\gamma}\log\left(\frac{\pi_{\theta}(y_{w}\mid x)}{1-\pi_{\theta}(y_{l}\mid x)}\right)}_{\text{Contrastive Log-"unlikelihood"}}\Bigg{]},$ (5)
@@
where $p_{w}=\sigma(z_{w}-z_{l})$ and quantifies the student policy’s confidence in correctly assigning the preference from $z_{w}=\log\pi_{\theta}(y_{w}\mid x)$ and $z_{l}=\log\pi_{\theta}(y_{l}\mid x)$, or the log-probabilities of the winning and losing responses, respectively. Hyperparameters $\gamma$ and $\alpha$ regulate the strength of the modulating term and weighting factor respectively. [^3] Note that we do not use $\alpha$ as it is traditionally used in focal loss for weighting class imbalances (Mukhoti et al., 2020). Instead, since the true preference distribution is unknown, we tune the empirical optimal value based on validation data and keep it fixed during training. As shown in Eq. 6, there is no shared parametrization between the policy and $\mathcal{O}$. $\mathcal{L}_{\mathcal{O}}$ is only a function of $\mathcal{O}$’s own parameters $\phi$ and the preference dataset, $\mathcal{D}_{\text{pref}}$. The Oracle parameters are not updated during policy training. Table 3 in the appendix provides the complete DRDO algorithm.
@@
Intuitively, DRDO’s first component in the loss effectively regresses over $\mathcal{O}$’s reward differences across all response pairs $(x,y_{1},y_{2})$ in $\mathcal{D}_{\text{pref}}$, using $\pi_{\theta}$’s own reward estimates, $\hat{r}_{1}$ and $\hat{r}_{2}$. Assuming $\mathcal{O}$’s reward estimates correctly reflect the true preference strength, this simple distillation component is precisely minimized when the student $\pi_{\theta}$ perfectly emulates $\mathcal{O}$’s reward difference assignment, without requiring it to explicitly match the pointwise rewards. The square over this term ensures that training is stabilized, since we allow this component to fully regularize the loss without applying any weighting terms.
@@
“Contrastive log-unlikelihood” in Eq. 5 is DRDO’s preference component. One can immediately draw a comparison with DPO which uses a fixed $\beta$ parameter (Eq. 3). DRDO uses a modulating focal-softened term, $(1-p_{w})^{\gamma}$ where $\pi_{\theta}$ learns from both deterministic and non-deterministic preferences, effectively blending reward alignment with preference signals to guide optimization. Intuitively, unlike DPO’s $\beta$ that is constant for every training sample, this modulating term amplifies gradient updates when preference signals are weak ($p_{w}\approx 0.5$) and tempering updates when they are strong ($p_{w}\approx 1$), thus ensuring robust learning across varying preference scenarios. When $\pi_{\theta}$ assigns high confidence to the winning response ($p_{w}\approx 1$, see Eq. 5), the focal loss contribution diminishes, reflecting minimal penalty due to strong deterministic preference signals. However, for harder cases with non-deterministic true preference ($p^{*}(y\succ y^{\prime})\approx(p^{*}(y^{\prime}\succ y)$), the focal term $\alpha(1-p_{w})^{\gamma}$ keeps DRDO gradient updates active and promotes learning even when preferences are ambiguous (Fig. 4 in Section B.6). This adaptive behavior ensures that DRDO maintains effective preference learning across varying preference strengths, where conventional methods like DPO struggle. See Lemma 1 and Lemma 2 for in-depth analysis. As shown in the gradient analysis in Appendix B.6 (Fig. 4) with empirical proof in Fig. 5 (bottom-right), this modulating term acts like DPO’s gradient scaling term, in that it scales the DRDO gradient when the model incorrectly assigns preferences to easier samples.
@@ 5 Experimental Setup
Our experimentals address two questions: How robust is DRDO alignment to nuanced or diverse preferences, in OOD settings?, and How well does DRDO achieve reward distillation with respect to model size? We empirically investigate these questions on two tasks: summarization and single-turn instruction following. Our experiments, including choice of datasets and models for each task are designed to to validate our approach as robustly as possible, subject to research budget constraints (see Section C.2 for more). We compare our approach with competitive baselines such as DPO (Rafailov et al., 2024b) and e-DPO (Fisch et al., 2024), including the supervised finetuned (baseline) versions depending on the experiment. Some minor notes on the experimental setup described below can be found in Appendix C.
@@
We evaluate this using the Reddit TL;DR summarization dataset (Völske et al., 2017; Stiennon et al., 2020) and CNN/Daily Mail Corpus (Nallapati et al., 2016). For robust out-of-domain (OOD) evaluation, we train models on Reddit TL;DR and use CNN/Daily Mail articles for an OOD test distribution. We split the original training data (denoted $\mathcal{D}_{all})$ based on human labeler confidence in their preference annotation ($h$igh-$c$onfidence vs. $\ell$ow-$c$onfidence) and token edit distance between the preferred and the dispreferred responses ($h$igh-$e$dit distance vs. $\ell$ow-edit distance). This results in two splits: $\mathcal{D}_{hc,he}$ and $\mathcal{D}_{\ell c,\ell e}$. Each contains $\sim$10k training samples, where the former comprises samples from the upper 50th percentile of the confidence and edit distance scores, mutatis mutandis. $\mathcal{D}_{hc,he}$ and $\mathcal{D}_{\ell c,\ell e}$ represent “easy” (deterministic) and “hard” (non-deterministic) preference samples where combined labeler confidence and string-dissimilarity act as proxy for the extreme ends of preference strengths/signals. See Appendix C.3 for more details. For each split, we compare our approach to DPO and e-DPO. All baselines are initialized with Phi-3-Mini-4K-Instruct weights (Abdin et al., 2024) with supervised fine-tuning (SFT) on Reddit TL;DR human-written summaries.
@@
We evaluate all baselines on the cleaned version of the Ultrafeedback dataset (Cui et al., 2024) This experiment is conducted on the OPT suite of models (Zhang et al., 2022), at 125M, 350M, 1.3B, and 2.7B parameter sizes. The student policy is trained with SFT on the chosen responses of the dataset, following Rafailov et al. (2024b). We exclude larger OPT models to focus on testing our distillation strategy with full-scale training, rather than parameter-efficient methods (PEFT; Houlsby et al. (2019); Hu et al. (2021)) to allow a full-fledged comparison considering all trainable parameters of the base model. For completeness and comparison across model families, we also include Phi-3-Mini-4K-Instruct following the same initialization.
@@
CNN/Daily Mail Corpus provides human-written reference summaries, so we use a high-capacity Judge to compute win-rates against baselines on 1,000 randomly-chosen samples. Following Rafailov et al. (2024b), we use GPT-4o to compare the conciseness and the quality of the DRDO summaries and baseline summaries, while grounding its ratings to the human written summary. See Appendix G for our prompt format. For instruction following on Ultrafeedback, we sample generations from DRDO and all baselines at various diversity-sampling temperatures and report win-rates on the Ultrafeedback test set against $\mathcal{O}$. Following Lambert et al. (2024), we consider a win to be when, for two generations $y_{1}$ and $y_{2}$, we get $r(x,y_{1})>r(x,y_{2})$, where $r(x,y_{1})$ and $r(x,y_{2})$ are the expected rewards (logits) from the policies being compared. For an additional OOD setting, we also evaluate DRDO (trained on Ultrafeedback) against baselines (all trained on Ultrafeedback) on AlpacaEval, which constitutes 805 instructions chosen to be representative of how human-preferred policies would respond. For this we evaluate using GPT-4 Turbo as suggested by the creators (Li et al., 2023). Hyperparameters and model configurations are given in Appendix E.
@@ 6 Results
Table 1 shows the win rates computed with GPT-4o as judge for 1,000 randomly selected prompts from the CNN/Daily Mail test corpus under OOD settings. We follow similar settings as Rafailov et al. (2024b) but further ground the prompt using human-written summaries as reference for GPT to conduct its evaluation (see Appendix G). For a fairer evaluation (Wang et al., 2024c; Goyal et al., 2023; Rafailov et al., 2024b), we swap positions of $\pi_{\theta}$-generated summaries $y$ in the prompt to eliminate any positional bias and evaluate the generated summaries on criteria like coherence, preciseness and conciseness, with the human written summaries explicitly in the prompt to guide evaluation. Note that the win rate does not necessarily equal $N/1000$ because for a small number of samples ($<$ 10) GPT-4o did not return a judgment. These samples were discarded from the denominator.
Table 1: Win rates computed for 1,000 randomly selected samples from the CNN/Daily Mail Corpus and all 805 samples from AlpacaEval for evaluation under OOD settings. This comparison evaluates the model against a baseline, presenting win counts alongside win rates. CNN/Daily Mail is evaluated with GPT-4o as judge and AlpacaEval, with GPT-4 Turbo as judge. $\mathcal{D}_{all}$, $\mathcal{D}_{hc,he}$, and $\mathcal{D}_{\ell c,\ell e}$ represent TL;DR training splits. “Gold” samples refer to human-written reference summaries used in prompts to ground the win rate computations.
Comparison (CNN/Daily Mail) # Wins Win Rate
@@
Comparison (AlpacaEval) # Wins Win Rate
@@
For all policies $\pi_{\theta}$ trained on all three splits of the training data—$\mathcal{D}_{all}$, $\mathcal{D}_{hc,he}$, and $\mathcal{D}_{\ell c,\ell e}$—we compute the win-rates of DRDO vs. e-DPO and DPO to evaluate how each method performs at various levels of preference types. Across all settings, DRDO policies significantly outperform the two baselines. For instance, DRDO’s average win rates are almost 79.4% and 79.9% against e-DPO and DPO, respectively. As we had hypothesized, DRDO-aligned $\pi_{\theta}$ is able to learn all preferences, deterministic and non-deterministic, effectively, and is superior at modeling $\mathcal{D}_{\ell c,\ell e}$ the subset containing more non-deterministic preference samples, without sacrificing performance on the deterministic subset ($\mathcal{D}_{hc,he}$) and overall ($\mathcal{D}_{all}$). This suggests that DRDO is more robust to OOD-settings at various levels of difficulty in learning human preferences.
Reward distillation
Fig. 2 shows results from our evaluation of DRDO’s reward model distillation framework compared to DPO and e-DPO as well as baseline SFT-trained policies on the Ultrafeedback evaluation data, when compared across various model parameter sizes and at varying levels of temperature sampling. We sample $\pi_{\theta}$-generated responses to instruction-prompts in the test set using top-$p$ (nucleus) sampling (Holtzman et al., 2019) of 0.8 at various temperatures $\in\{0.2,0.5,0.7,0.9\}$.
Figure 2: Average Ultrafeedback win-rates computed with DRDO’s Oracle reward model against all baselines—SFT, DPO and e-DPO—at various diversity sampling temperatures ($T$).
DRDO significantly outperforms competing baselines, especially for larger models in the OPT family. DRDO-trained OPT-1.3B, OPT-2.7B, and Phi-3-Mini-4K-Instruct achieve average win rates of 76%, 74%, and 72%, respectively, across all baselines. This is notable as responses are sampled on unseen prompts, and DRDO’s policy alignment is reference-free. DRDO’s robustness to diversity sampling further boosts performance, up to an 88% win rate against DPO with the Phi-3-Mini-4K-Instruct model. At lower temperatures, DRDO’s posted gains are more modest. Our results also indicate that performance is correlated with model size, as DRDO policies of the same size as the Oracle (1.3B) show the strongest gains. In smaller models, results are more mixed. [^4] This may be due to spurious length-correlation (Singhal et al., 2024; Meng et al., 2024), as smaller models generated avg. $\sim$211.9 tokens compared to their policy counterparts’ $\sim$190.8 tokens, which is closer to the original preferred responses in Ultrafeedback ($\sim$168.8). See Appendix E for more details. DRDO shows moderate improvement and posts smaller gains against SFT models.
@@
7 Analysis
Table 2 shows example generations from DRDO and a competitor where the DRDO example was preferred by the automatic judge. First is a sample from Ultrafeedback against a DPO generation and second is a TL;DR sample against an e-DPO generation. We see that the DRDO responses are more concise and on-topic while the competitor output condescends to the user, or includes extraneous text about fulfilling the request.
Using GPT-4o as a judge to approximate true human preferences may be prone to bias, so we further validate our approach by investigating Oracle reward advantage over the above mentioned human written summaries as well as on-policy generations from an SFT-trained model. Fig. 3 shows the computed expected reward advantages on the CNN Daily TL;DR evaluation set, sampled according to the method outlined in Sec. 6. Rewards were computed using our Oracle (trained with Phi-3) on these sampled generations and normalized. To compute the advantage, we used human written summaries (Fig. 3). DRDO improves performance across various temperature samplings over a baseline SFT policy, and brings in a considerable performance gain over competitive baselines like DPO and ensemble-based e-DPO while also being robust to OOD settings.
Figure 3: Oracle expected reward advantage on CNN/Daily Mail articles.
Our evaluations vs. “gold” summaries in Table 1 also demonstrate where bias may arise in the GPT-4o evaluation (win counts and win rates are shown for gold samples). GPT-4o narrowly prefers generated summaries to human-written ones. One possible reason could be that Reddit TL;DR data is massive and crowd-sourced which naturally results in noisy labels. However, under this experiment too, DRDO-trained policies are better-performing than e-DPO and DPO, by about 1.5–6%. We should note that human summaries may contain more implicit diversity than generated summaries, and this may demonstrate the “regression to the mean” effect in LLM generation (Wu et al., 2024).
Our results on AlpacaEval demonstrates DRDO’s instruction following ability in OOD settings. Table 1 shows that, within controlled evaluation settings, GPT 4-Turbo consistently prefers DRDO responses 62.03% and 61.61% times compared to e-DPO and DPO respectively.
Further, we conducted an ablation on 40 randomly sampled prompts from the CNN/Daily Mail test set, using GPT-4o as a judge, comparing a full DRDO-aligned model (Phi-3-Mini-4K-Instruct aligned using DRDO policies trained on Reddit TL;DR) to DRDO with only reward distillation and only contrastive log-unlikelihood. In this subsample, DRDO won over DRDO with only contrastive log-likelihood 85%-15%, and over DRDO with only reward distillation 90%-10%. This shows that both components are critical to DRDO’s success. In similar settings, we also examined the sensitivity of DRDO’s $\gamma$ vs. DPO’s $\beta$ on randomly sampled prompts from the Ultrafeedback evaluation set. We find that despite DRDO’s exclusion of the KL constraint on $\pi_{\text{{ref}}}$, DRDO is still able to recover more expected rewards signifying a more optimal trade-off between reward optimization and divergence from $\pi_{\text{ref}}$. At $\gamma=2$, DRDO wins on 5% more samples than DPO despite a slightly larger KL-divergence. This suggests that explicit reward distillation in DRDO with preferences being learned adaptively (via $\gamma$) makes it more Pareto-optimal especially in the presence of non-deterministic preferences. See Appendices F.1 and J for a more comprehensive discussion.